OpenStudy (anonymous):

An unconditional acceptance into a graduate program at a university will be given to students whose GMAT score plus 400 times the undergraduate grade point average is at least 1800. Robbin's GMAT score was 800. What must her grade point average be in order to be unconditionally accepted into the program?

OpenStudy (anonymous):

(GMAT score + 400) X (Undergrad Grade Point Average) >= 1800 \[1200 \times Undergrad Grade Point Average\ge1800\] \[Average \ge 1800/1200 = 3/2=1.5\] Bigger or equal to 1.5

OpenStudy (anonymous):

so her GPA must be at least 1.5?

OpenStudy (anonymous):

Wait, I might've misunderstood your question. "GMAT score plus 400 times the undergraduate grade point average" might mean GMAT Score + 400*GPA If it is like this, then \[800 + 400 \times GPA \ge 1800\] so \[400 \times GPA\ge1000\] so \[\times GPA\ge1000/400=2.5\] So 2.5 or higher.

OpenStudy (anonymous):

ok that's easier to understand.. thanks

OpenStudy (anonymous):

Welcome^^